OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL AND WILL FAN!!! After a discount of 10%, the price of a jacket, including tax, is $70.79. The following equation can be used to find x, the price of the jacket without the discount: x - 0.1x + 0.06(x - 0.1x) = 70.79 What is the price of the jacket without the discount? A. $71.74 B. $74.20 C. $75.31 D. $82.31

OpenStudy (anonymous):

x = 79.18

OpenStudy (anonymous):

The answer is B Let's try to solve the equation. we have: x-0.1x +0.06(x-0.1x)=70.79 (x-0.1x)(1+0.06)=70.79 (by factorizing x-0.1x) x(1-0.1)(1+0.06)=70.79 (by factorizing x) x(0.9)(1.06)=70.79 0.954x=70.79 0.954x/0.954=70.79/0.954 x=74.20 I hope I have helped you.

OpenStudy (anonymous):

@Jaspharmo thanks you really did help!!! :)

OpenStudy (anonymous):

You're welcome! It was a pleasure to try to help. And thank you sooo much for the medal!